A grasshopper can jump maximum distance of 1.6.it spends negligible time on ground.How far can it go in 10 seconds

Grade:11

1 Answers

Arun
6 years ago
grasshopper jumps maximum distance when Ф = 45 deg

Horizontal component of velocity = u /√2 

R = U² / g         u = √gR  = √10*1.6  = 4 m/sec

Distance in 10 sec:  4/√2 * 10  = 20√2 m